# LookAhead: Preventing DeFi Attacks via Unveiling Adversarial Contracts

### v4

@@ 6.2.3. Comparison with the Baseline
Our meta classifier pushes the boundary even further by combining the capabilities of transformer model with the capabilities of other candidate classifiers to achieve F1-Score as high as $0.8966$. Furthermore, as described in §5.1, contract verification status feature introduces some bias into our dataset, to provide a fair comparison, we also evaluate the performance of our models with $verified$ feature removed. Experiments show that even with the feature $verified$ removed from our training and evaluation dataset, our classifiers still significantly outperform the most recent model published by Forta (released on February 6, 2023), achieving an F1-Socre of 0.8696, highlighting the superiority of LookAhead’s solution.
6.3. RQ2: Efficiency
Efficiency testing evaluates LookAhead’s resource consumption during its operating time, such as processing power requirements, memory usage, and detection throughput. All these contribute to delays in the identification of attackers and potential victims. Given the frequency of smart contract deployment transactions in decentralized finance, LookAhead should be able to quickly respond and react to potential attack activities within typical rescue time frames.
The most time-consuming step in LookAhead’s detection pipeline is the the contruction of PSCFT. Thanks to the rapid advancement of ML, classifier inference can be done on either CPUs or GPUs in a batched manner, which means in most cases it is a significantly faster process than binary lifting and analysis. Additionally, for external data augmentation pipelines that require sending API requests to outside sources, we can often easily employ multi-threading techniques to achieve high data retrieval concurrency. Thus, it can be assumed that the time required for analyzing a smart contract is roughly equivalent to the time needed for constructing its PSCFT.
To evaluate the efficiency, we measured the analysis time for all adversary contracts. The results showed that, except for one large contract with bytecode close to the maximum size, which took about $455$ seconds, most other contracts’ PSCFTs were constructed in under $20$ seconds, with an average time of $7.34$ seconds. Additionally, by considering the rescue time frames (cf. §3.2.4) of their corresponding DeFi attacks, we found that in $92.8\%$ of the cases, our system could respond in time by detecting the adversary contract before the adversary transaction occurred, leaving plenty of room for defense actions to be taken by affected parties and proving the viability of the LookAhead system.
6.4. RQ3: Practicality
We consider the practicality of the LookAhead system crucial to the validity of our research, therefore we conducted experiments measuring the detection performance of LookAhead to verify its effectiveness in realistic, dynamic environments.
From August 14 to August 18, 2024, we deployed the LookAhead system to a cloud server and conducted a live experiment to measure LookAhead’s performance in real-world scenarios. We use a meticulous review process that require us to 1) Wait for $2$ weeks (September 1) to ensure attacks have been confirmed, 2) Manually review the potential adversarial contracts identified by LookAhead, 3) Validate the detection results by inspecting follow-up transactions involving the adversarial contracts, 4) Calculate the total amount of financial damage caused by valid DeFi attacks, 5) Measure the time gap between the deployment of an adversarial contract and the identification of the contract by LookAhead.
@@ 6.4. RQ3: Practicality
Out of $39$ adversarial contracts flagged by LookAhead, $23$ involved follow-up transactions, in which we found $10$ true positives after going through the manual investigation process that were used by real attackers to perform DeFi attacks. These attacks caused over $50,000$ USD in financial losses. In terms of detection efficiency, except only one case where the deployment and attack transaction were confirmed within the same block, LookAhead produced a response for all the attack events before the attacker actually initiated the adversarial transaction. Additionally, after thorough analysis, we observed that almost all the false positives are related to arbitrage, showing that the behaviour exhibited by arbitrage trading resembles some patterns in adversarial activities. For example, the dependency on flashloans, the high number of token ralated function calls, etc.
@@ 7. Discussion
Threats to Validity. Limited by the API responses given by the blockchain explorers, we can only go as far as pinpointing the date of contract verification without knowing the exact time. Thus we cannot determine whether verification was done at the time of deployment or not. Additionally, due to the scarcity of actual adversarial contract samples, the ratio between the two types of samples in our final dataset is approximately $1:40$. This imbalance could affect the performance of our models. Furthermore, benign contracts were selected based on empirical observations and the assumption that adversarial contracts typically have fewer unique interactions with other accounts. There is still a chance that we mistakenly treated undisclosed adversarial contracts as benign contracts during training and testing, which could have an impact on the validity of our results. Smart contracts can be created through internal transactions triggered during the execution of a normal transaction. Those contracts are not included in our dataset.
Limitations. Our current method has certain limitations in terms of the spectrum of adversarial contracts that LookAhead is capable of detecting. When selecting features, we rely on existing attack patterns seen in DeFi attacks involving financial losses and those in prevalent attack scenarios. Consequently, there is a probability that we might have missed attacks that do not result in financial damages and certain adversarial contracts with novel attack vectors. Furthermore, in the cat-and-mouse game of DeFi attack and defense, attackers may evade our detection by using techniques like code obfuscation to alter the CFG of contract programs or by utilizing KYC-required exchanges as funding sources that makes it harder for our system to recognize adversarial intentions.
@@ 8. Related Work
There have been many studies concerning the detection of abnormal contracts, primarily targeting Ponzi schemes and honeypots. They share similarities in methods which could be categorized into code similarity detection (Bartoletti et al., 2020; Sun et al., 2020), machine learning (Chen et al., 2020; Hu et al., 2022; Ji et al., 2023; Zheng et al., 2023), and symbolic execution (Chen et al., 2021). However, to the best of our knowledge, this is the first academic work that focuses on identifying adversarial contracts, and the first to support detection across all types of DeFi attacks. Forta (For, 2024) also proposed a method that uses machine learning techniques to identify adversarial contracts, but their approach is limited to bytecode-level analysis. Aside from Forta, the only similar work is BlockWatchDog proposed by Yang et al. (Yang et al., 2024), which uses cross-contract static dataflow analysis technique to detect adversarial contracts in order to identify reentrancy vulnerabilities. However, their focus is solely on reentrancy attacks, which account for only $6.7\%$ of the attacks in our dataset, and their method relies heavily on knowing the victim addresses in advance.
@@ 9. Conclusion
In this paper, we propose a new direction for detecting DeFi attacks, i.e., detecting adversarial contracts instead of adversarial transactions. We build the first dataset for training PSCFT-based transformers, classifiers and meta classifiers for the binary classification of adversarial contracts. Based on this dataset, we found that the use of private mempool services by DeFi attackers is becoming increasingly common, rendering traditional transaction-based detection methods ineffective in most cases. Through the analysis of DeFi attack patterns, we extract a multi-dimensional feature set along with a special tokenization technique that is able to capture the intrinsic behaviours hidden in adversarial contracts. Utilizing these features, we build a transformer, multiple candidate classifiers and a meta classifier and evaluate them comprehensively. Experiments show that our method performs exceptionally well in detecting and preventing DeFi attacks with our models reaching F1-Scores as high as $0.8966$ and a false positive rate as low as $0.16\%$.

### v6

@@ 6.2.3. Comparison with Previous Works
Forta is an NLP-based method that relies solely on EVM bytecode, while our method is based on a combination of multi-dimensional features and semantic tokenization (PSCFT) of smart contracts. We perform a direct comparison between their model and our transformer model first, then we measure how much benefits can our meta classifier approach provide over Forta’s implementation. As the results in Table 1 indicate, our transformer alone can achieve much higher F1-Score than Forta, demonstrating the advantage of our PSCFT design. LookAhead’s meta classifier pushes the boundary even further by combining the capabilities of transformer model with the capabilities of other candidate classifiers to achieve an F1-Score of $0.8966$. As described in §5.1, contract verification status feature introduces some bias into our dataset, to provide a fair comparison, we also evaluate our models with $verified$ feature removed. The results show that even without this feature, our classifiers still significantly outperform the most recent model published by Forta (released on February 6, 2023), achieving an F1-Socre of $0.8696$, highlighting the effectiveness of our solution.
6.2.4. Time Series Cross-Validation
As outlined in 6.1.1, our evaluation respects the chronological order of contract deployment, accounting for both the evolving nature of adversarial behaviours and LookAhead’s design to detect future attacks based on historical data. Table 1 shows the model performance results obtained by using the entire dataset with a 4:1 chronological split. We identify XGBoost as the optimal candidate classifier and KNN as the best-performing meta model. To validate the most appropriate candidate and meta models for our system and ensure its robustness under varying data conditions, we conduct evaluations using expanding window cross-validation. During this process, the dataset is chronologically divided into five equal splits. The training window starts with the first split and progressively expands by including additional splits, while the subsequent split serves as the test set in each iteration.
After cross-validation, we conclude that XGBoost achieves the highest average F1-score (0.8173), demonstrating stable performance across varying data splits. While it did not always outperform others in every split, its overall consistency reaffirms its suitability as the optimal candidate classifier for our system. For the meta classifiers, which combine the optimal candidate model XGBoost with the transformer, the KNN-based model achieves the highest average F1-score ($0.8355$). However, its advantage over other models were marginal, suggesting that the choice of meta classifier models has limited impact on overall performance.
6.3. RQ2: Interpretability
6.3.1. Empirical Features
Understanding the weight distribution of features is useful for gaining insights into the underlying decision-making process of the models. To this end, we apply the interpretability method SHAP (Shapley Additive exPlanations) (Lundberg and Lee, 2017) to analyze the importance of empirical features for each candidate model.
The results reveal that both deployment and implementation features play significant roles in guiding the models to a decision. In the best-performing XGBoost model, the top three most important features are avg_token_call_count, verified and fund_from_Anonymous, all of which also ranked within the top 10 across other candidate models. Additionally, balanceOf_call_count, fund_from_Safe and func_count consistently appear within the top 15 in at least three candidate models, indicating their stable contribution to the model’s decision-making process. The notable importance of these features may be attributed to their strong correlation with typical attack patterns, allowing the model to effectively capture behavioural characteristics of adversarial contracts.
6.3.2. PSCFT
The design of PSCFT abstracts the IR of a contract into a more concise and semantically meaningful format. To concretely demonstrate how this design enhances model performance, we compare the transformer classifier trained on PSCFT with one trained on the raw IR generated by Gigahorse. The results show a significant performance gap, where the model using raw IR achieves an F1-score of only 0.4068, substantially lower than the 0.7222 achieved by the PSCFT-based model.
This performance discrepancy highlights the limitations of directly using raw IR, which contains numerous low-level instructions (e.g., arithmetic operations, memory reads/writes) and excessive variables that obscure critical call patterns and their control flows. Such extraneous details increase input complexity, making it harder for the model to identify meaningful behavioural patterns relevant to adversarial contract. Our pruning process can effectively remove redundant instructions and enable semantic recovery, allowing PSCFT to focus on critical external calls and control flows, leading to clearer and more reliable model decisions and improved prediction accuracy.
6.4. RQ3: Efficiency
LookAhead  aims to detect adversarial contracts and enable rapid responses within the rescue time frame (see §3.2.4). To this end, we evaluate the efficiency using all adversarial samples.
We employed a linear detection pipeline, which processes each sample through the following steps: 1) retrieve deployment features via external APIs, 2) execute our custom-designed Souffle Datalog and PSCFT generation scripts to extract implementation features, and 3) preprocess the data and feed it into pre-trained XGBoost and transformer classifiers. Their outputs are then combined by a pre-trained meta-classifier for the final prediction. We measure the time taken at each step.
The results showed that, on average, external API requests took $3.78$ seconds, while data preprocessing and model inference required only $0.01$ seconds. The most time-consuming step is binary lifting and analysis. Except for one large contract (taking $458$ seconds), most PSCFTs were generated under $20$ seconds, averaging $7.52$ seconds. Furthermore, by considering the rescue time frames of the corresponding DeFi attacks, we found that in $92.8\%$ of the cases, our system successfully detected the adversarial contract before the adversarial transaction occurred. Notably, efficiency can be further improved by employing multi-threading and local blockchain nodes, which could substantially reduce the time required to obtain deployment features.
6.5. RQ4: Practicality
Setup We conducted live experiments targeting the Ethereum mainnet across two time periods, 14–18 August 2024 and 23 January–9 February 2025, by deploying the LookAhead system to the server. The system continuously monitored the transactions in each new block via RPC providers (Alc, 2024) in near real-time. Upon detecting a contract creation (through a normal transaction), scripts were executed to extract features, which were then preprocessed and fed into the pre-trained classifiers (XGBoost and KNN, the best-performing setup as identified in RQ1). If classified as adversarial, the contract address was recorded for further verification.
Review Process We use a meticulous review process as follows to verify the outputs:
  1. 1)
    Wait for $2$ weeks after each experimental period to ensure attacks have been confirmed;
  2. 2)
    Manually inspect follow-up transactions involving the contracts to validate adversarial nature;
  3. 3)
    Analyze and record key data, e.g. the financial damage caused by confirmed attacks, the time window between contract deployment and corresponding adversarial transaction.
Table 2. Representative true positive adversarial contracts detected during the live experiment, involved in attacks triggered by different root causes. Window indicates the time interval between contract deployment and the adversarial transaction. Runtime denotes the detection time, and Victim specifies whether the victim address was exposed during contract creation.
Address Time Window Runtime Victim Root Cause Loss (\$)
0x9f2790 2024/8/16 25344s 11.53s No Price Manipulation 15.5k
0x90744c 2024/8/16 636s 4.84s Yes Stale Oracle 21.5k
0x152619 2024/8/18 336s 5.67s Yes Access Control Flaw 1.2k
0x172133 2025/1/23 3504s 6.12s No Absence of Input Validation 2.5k
0x203f20 2025/2/6 120s 14.35s Yes Arbitrary Transfer 37k
0xbfe102 2025/2/6 96s 6.82s Yes Reentrancy 8.7k
Results During the experiment period, LookAhead analyzed a total of $23,463$ contracts and flagged $129$ as adversarial. Among these, $62$ involved follow-up transactions, in which we finally found $26$ true positives that were used to perform attacks, resulting in over \$150,000 USD in financial losses. These attacks were driven by various underlying issues, with Arbitrary Transfer (Ma et al., 2023) being the most common, involving 6 contracts. Table 2 presents representative adversarial contracts involved in attacks stemming from different root causes, highlighting the universality of our approach, which effectively identifies various types of adversarial behaviours.
@@ 6.5. RQ4: Practicality
In terms of detection efficiency, except only two cases where the deployment and attack transaction were confirmed within the same block, LookAhead produced a response for all the attack events before the attacker actually initiated the adversarial transaction. Additionally, we observed most of the false positives are related to arbitrage, showing that arbitrage behaviour resembles some patterns in adversarial activities, such as reliance on flashloans and frequent token-related function calls. To evaluate potential false negatives, we reviewed the alerts posted by security company (e.g. (blo, 2024; cer, 2024; slo, 2024)) during the experiment period, as manually checking for false negatives among a large number of contracts is challenging. This process identified four reported attacks that fell within our detection scope (§3.2.1), all of which had already been correctly flagged.
The fact that, during our live experiment, LookAhead was able to respond in a timely manner before the majority of real-world attacks occurred, and successfully detected all externally reported attacks, further demonstrates its practicality in real-world scenarios.
@@ 7. Discussion
7.1. Threats to Validity
1) Limited by the APIs given by the blockchain explorers, we cannot determine whether contract verification occurred upon deployment; 2) Our dataset suffers from a class imbalance of approximately $1:40$ due to the scarcity of adversarial contracts, which may affect model performance even after oversampling; 3) Benign contracts were selected based on heuristic approach. There is still a possibility that we have mislabelled some adversarial contracts as benign ones; 4) As discussed in §4.1.2, our study does not consider contracts deployed via internal transactions. The inherent differences between the two deployment methods may affect certain deployment features, introducing potential bias to our classifier when detecting contracts deployed via internal transactions.
7.2. Limitations
In addition to the inherent limitations of our threat model definition (§3.2), there are mainly two aspects of LookAhead’s design that could potentially hinder its performance.
Detection Evasion Since our classifier relies on extracted features and historical attack patterns, it remains susceptible to evasion techniques. In the cat-and-mouse game of DeFi attack and defense, attackers may adapt their strategies to bypass our feature-based and model-driven approach. For instance, they could adopt code obfuscation to alter the contract’s CFG or utilize KYC-required exchanges as funding sources, making it harder for our system to recognize adversarial intentions.
NLP-based Approach LookAhead’s NLP-based models may not make full use of the control flow graph structure present in smart contract code. More advanced ML architectures such as Graph Neural Network (GNN) could be used in future works in combination with a more intricate representation design that capture both intra-function CFGs and inter-function call relationships.
@@ 8. Related Work
Adversarial Transaction Detection. Prior research primarily focus on detecting adversarial transactions used in DeFi attacks. Some works (Chen et al., 2020; Grossman et al., 2017; Rodler et al., 2018; Wu et al., 2022) modify EVM to obtain more detailed transaction execution information for detecting adversarial transactions. BlockGPT (Gai et al., 2023) applies large language model to detect abnormal transactions. Qin et al. (Qin et al., 2023b) introduce the Execution Property Graph to enable timely detection. However, these methods become ineffective in identifying attacks when attackers leverage private mempool services to initiate adversarial transactions.
Adversarial Contract Detection. The study of adversarial contract detection in DeFi attacks remains relatively underexplored. Forta (For, 2024) adopts an NLP-based approach for adversarial contract detection; however, its reliance on simple bytecode-level analysis results in limited effectiveness. Beyond Forta, the only comparable work is BlockWatchDog (Yang et al., 2024). However, their approach is restricted to reentrancy attacks and relies on prior knowledge of victim addresses, thereby limiting its practical utility in real-world scenarios. Our method integrates multi-faceted features with semantic tokenization to allow more effective detection across various DeFi attack types.
@@ 9. Conclusion
In this paper, we propose a new framework for effectively detecting DeFi attacks. We build the first large-scale comprehensive dataset for training classifiers for the binary classification of adversarial contracts. Through the analysis of DeFi attack patterns, we extract a multi-dimensional feature set along with a special tokenization technique that is able to capture the intrinsic behaviours hidden in adversarial contracts. Utilizing these features, we carefully design multiple classifiers and evaluate them comprehensively. Experiments show that our method performs exceptionally well in detecting adversarial contracts compared to the previous state-of-the-arts with our models reaching F1-Scores as high as $0.8966$ and a false positive rate as low as $0.16\%$ on our labeled datasets.
10. Data Availability
To promote open science, we have made our dataset and source code publicly available (loo, 2024).
Acknowledgements.
This work is sponsored in part by National Key Research and Development Program of China (2023YFB2704000).